a man is 2years more than thrice his son age,if the sum of their age is 46 how old are they?
Answer by ankor@dixie-net.com(22740)

: a man is 2years more than thrice his son age,
m = 3s + 2
:
if the sum of their age is 46
m + s = 46
m = (46 - s), this form for substitution
:
Substitute (46-s) for m in the 1st equation
46 - s = 3s + 2
46 - 2 = 3s + s
44 = 4s
s = 11 yrs is son's age
then
46 - 11 = 35 is man's age
:
:
Check solutions in the statement;
": a man is 2 years more than thrice his son age,"
35 = 3(11) + 2